Part B How many grams of NH, can be produced from 2 84 mol of N2 and excess H. Express your answer numerically in grams. View Available Hint(s) g NH Submit Part C How many grams of Ha are needed to produce 11.31 g of NHs? Express your answer numerically in grams. View Available Hints) g Ha Submit
Answer #1

part B

N2 + 3H2 ----> 2NH3

1 mol N2 = 3 mol H2

No of mol of N2 = 2.84 mol

no of mol of NH3 produced = 2.84*2 = 5.68 mol

amount of NH3 produced = 5.68*17 = 96.56 g

part C

N2 + 3H2 ----> 2NH3

1 mol N2 = 3 mol H2

no of mol of NH3 = 11.31/17 = 0.665 mol

no of mol of H2 = 0.665*1/2 = 0.3325 mol

mass of H2 required = 0.3325*2 = 0.665 g